OpenStudy (anonymous):

Which of the following equations is equivalent to y = lnx? A. x = y^e B. y = x^e C. y = e^y D. y = e^x

OpenStudy (solomonzelman):

\(\large\ln(x)\) is same as \(\large\log_{e}(x)\)

OpenStudy (anonymous):

Ok ok, so would I take the 'e' and sort of "bump" it over to the 'y' slot making e^y?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\ln(x) }\) \(\large\color{black}{ \displaystyle y=\log_{e}(x) }\) then apply the rule: \(\huge\color{black}{ \displaystyle \color{red}{\rm a}=\log_{\color{green}{\rm b}}(\color{blue}{\rm c})~~~~~\Longrightarrow ~~~~~ \displaystyle \color{green}{\rm b}^{\color{red}{\rm a}}=\color{blue}{\rm c} }\)

OpenStudy (solomonzelman):

this rule that I drew with many colors, is what you should use for your final answer.

OpenStudy (solomonzelman):

actually, the way you wrote the options, none of them are correct-;(

OpenStudy (anonymous):

Oh no ;(. Really?

OpenStudy (solomonzelman):

you made a typo in your options. check them

OpenStudy (solomonzelman):

but, do it without options and tell me what you get

OpenStudy (anonymous):

I would first set it as \[y = \log_{e} x\] and then turn it into \[e ^{y} = x\]

OpenStudy (anonymous):

Oh and option C should be x=e^y. Sorry

OpenStudy (anonymous):

Was that right?

OpenStudy (solomonzelman):

yes, that was right

OpenStudy (solomonzelman):

well done

OpenStudy (anonymous):

Yay! Thank you so much!!!